### v1

@@ I Introduction
The dynamic nature of dark energy is one of the alternatives to the cosmological constant which can help us to alleviate the current tensions. Modified theories of gravity are successful in giving a dynamical nature to the dark energy and dark matter phenomenon. For example, the quintessence model is introduced to explain the late-time cosmic expansion with a scalar field, instead of trivial cosmological constant $\Lambda$ [11] [^1] There are also other models that can mimic the dark energy phenomenon, with out introducing any additional degrees of freedom [12, 13, 14, 15, 16, 17, 18].. Later this model was generalised to more general theories of scalar-tensor gravity, like $k-$essence [19], Horndeski gravity [20], beyond Horndeski [21] and DHOST theories [22]. Vector-tensor theory is also introduced to explain the dynamical dark energy, for example [23]. Vector-tensor dark energy was also generalised to a Generalised Proca theory [24], further extended to Proca-Nuevo [25]. The cosmology of vector-tensor theories are studied in [26, 27, 28]. The dark energy models with scalar field and the vector field were extended to interaction with cold dark matter. In the literature, the interaction is usually introduced with phenomenological approach. In [29] a general Lagrangian for the interacting scalar field dark energy and dark matter is introduced, in which an interaction term is $\partial_{\mu}\phi u^{\mu}$, where $u^{\mu}$ is the four velocity of the cold dark matter. This interaction is then further extended to generalised scalar-tensor theory [30]. For the dark energy modelled by vector field $A_{\mu}$, a similar interaction term, $Z=A_{\mu}u^{\mu}$ was introduced in [31]. Using similar methods of [29], a generalised interacting Lagrangian with momentum transfer is studied [32], where new interaction terms were found.
@@
In the work, we propose an interacting dark sector, in which the interaction term between CDM and dark energy is $Z\equiv u^{\mu}\partial_{\mu}\phi$. We introduce this term into the Lagrangian by $L_{\rm int}=G_{3}(\phi,X,Z)\nabla_{\mu}\partial^{\mu}\phi$. This is a novel construction, motivated from the fact that in the Horndeski theory, there exist higher derivative terms which is expected to modify the effective gravitational coupling at the level of linear perturbation. In this work, we study the effect of adding an interaction in the $G_{3}$ term of the Horndeski theory.
@@ II The theory
where $u_{{\rm c}}^{\mu}$ is the four-velocity of the cold dark matter component in the universe. The coefficient $\alpha_{1}$ is a dimensionless coefficient. The matter action for radiation and baryon is represented by $S_{m}$ and given by Schutz-Sorkin action
@@ IV.2 Scalar Modes
$\left[2M_{\rm P}^{2}H+F_{,Z}\dot{\phi}^{2}-G_{,X}\dot{\phi}^{3}\right]\alpha=2M_{\rm P}^{2}\zeta-n_{\rm m}\delta l_{\rm m}-n_{\rm c}\delta l_{\rm c}\,.$ (25)
@@ VIII.1 $H(z)$ data
VIII.2 RSD data
@@
The set of orthogonal functions (grammar) that we have used in our GA analysis are $x$ and $x^{x}$. We have chosen the crossover rate and the mutation rate as $0.75$ and $0.25$ respectively, with $1500$ generations. The algorithm will be terminated once the $\chi^{2}$ does not improve for a few hundred generations. The resulting curve is given in the figure 5. We also plotted the $f\sigma_{8}$ predicted by the model choosing one particular value of $\gamma=0.01$, which is the new parameter in the model. It is clear that the value of $f\sigma_{8}$ in the redshift close to today deviates from the $1\sigma$ predicted by GA, using the RSD data compilation as given in the table 2.
@@
In fact, we have only analysed this model with a particular value of the new parameter. It would be interesting to know if there are other values for the $\gamma$, which can fit with the data. To study this we need to do a complete parameter estimation using MCMC methods. This will also allow us to confront the model with other data sets like cosmic microwave background and other large scale structure data, which will be done in a future work.
@@ IX Conclusion
There are several future directions that are interesting to pursue, gaining intuitions from the current work. One direction is to generalize the concrete model to a general tracker solution with $\dot{\phi}H^{p}=\text{constant}$ as explored in [46], and following similar analysis and see if the new parameter $p$ can help in reducing the $f\sigma_{8}$ in to $1\sigma$ region of the genetic algorithm pipeline, using the same data. Another direction is to generalise the interacting model itself as studied in the context of Horndeski theory [73]. Then it would be interesting to see the behavior of the $f\sigma_{8}$. Even though, we did not see any additional propagating modes on the perturbation around the FLRW background, it would be interesting to see if any spurious modes are present in this model. One approach is to do a perturbation analysis on the Bianchi type I spacetime as done in [74, 75, 76]. Another obvious direction is to study the constraints of this model by implementing the Einstein - Boltzmann code and confront with the cosmological data. It is also interesting to study the constraints of this new model with integrated Sachs-Wolfe effect and the large scale structure cross correlation data [77, 78], which is a powerful tool to constraint the dark energy models.
@@ Appendix B Behaviour of $G_{cc}$ and $G_{bb}$ along the redshift
In this appendix we show the behaviour of the $G_{cc}/G_{N}$ and $G_{bb}/G_{N}$ along the variable $\mathcal{N}=\ln(a)$. As we can see that there is enhancement in the ratio between $G_{cc}(G_{bb})$ and $G_{N}$ in the dark energy domination era.

### v2

@@ I Introduction
The dynamic nature of dark energy is one of the alternatives to the cosmological constant which can help us to alleviate the current tensions. Modified theories of gravity are successful in giving a dynamical nature to the dark energy and dark matter phenomenon. For example, the quintessence model is introduced to explain the late-time cosmic expansion with a scalar field, instead of trivial cosmological constant $\Lambda$ [11] [^1] There are also other models that can mimic the dark energy phenomenon, with out introducing any additional degrees of freedom [12, 13, 14, 15, 16, 17, 18].. Later this model was generalised to more general theories of scalar-tensor gravity, like $k-$essence [19], Horndeski gravity [20], beyond Horndeski [21] and Degenerate Higher Order Scalar Tensor (DHOST) theories [22]. Vector-tensor theory is also introduced to explain the dynamical dark energy, for example [23, 24]. Vector-tensor dark energy was also generalised to a Generalised Proca theory [25], further extended to Proca-Nuevo [26]. The cosmology of vector-tensor theories are studied in [27, 28, 29]. The dark energy models with scalar field and the vector field were extended to interaction with cold dark matter. In the literature, the interaction is usually introduced with phenomenological approach. In [30] a general Lagrangian for the interacting scalar field dark energy and dark matter is introduced, in which an interaction term is $\partial_{\mu}\phi u^{\mu}$, where $u^{\mu}$ is the four velocity of the cold dark matter. This interaction is then further extended to generalised scalar-tensor theory [31]. For the dark energy modelled by vector field $A_{\mu}$, a similar interaction term, $Z=A_{\mu}u^{\mu}$ was introduced in [32]. Using similar methods of [30], a generalised interacting Lagrangian with momentum transfer is studied [33], where new interaction terms were found.
@@
In the work, we propose an interacting dark sector, in which the interaction term between CDM and dark energy is $Z\equiv u^{\mu}\partial_{\mu}\phi$. We introduce this term into the Lagrangian by $L_{\rm int}=G_{3}(\phi,X,Z)\nabla_{\mu}\partial^{\mu}\phi$. There are a few motivations for introducing this new interaction model. First, while constructing an interacting scalar dark energy the general Lagrangian is given as $L(n^{\mu},u^{\mu},\partial_{\mu}\phi,\phi)$, where $n^{\mu}$ is the density current. However, the exploration of the cosmological phenomenology is restricted to assuming a quintessence scalar field together with interaction term $f(Z)$, or the Horndeski theory with the same interaction term. However, it is possible to have interaction term with higher order derivative terms, which can give non-trivial phenomenology. Second, it is already known that in the Horndeski theory, there exist higher derivative terms which modify the effective gravitational coupling at the level of linear perturbation. In the $G_{3}$ Horndeski theory for some value of the parameters, the Integrated Sachs-Wolfe (ISW)-Galaxy cross correlation is found to have negative values for some regions of parameters [37, 38]. It is noticed that, while lowering the value of the $G_{{\rm eff}}$ the correlation turns to positive, but higher values of the $G_{\rm eff}$ can have better fit with the observations. Considering this fact, it is interesting to introduce features like an interaction term in the $G_{3}$ function of the Hodrneski theory to see, if the negative cross-correlations can be removed. However, as a first step, we need to make sure that this introduction does not introduce any instabilities.
@@ II The theory
where $u_{{\rm c}}^{\mu}$ is the four-velocity of the cold dark matter component in the universe. The coefficient $\alpha_{1}$ is a dimensionless coefficient. The matter action for radiation and baryon is represented by $S_{m}$ and given by Schutz-Sorkin action [50, 51, 52, 53],
@@ IV.2 Scalar Modes
$\left[2M_{\rm P}^{2}H+F_{,Z}\dot{\phi}^{2}-G_{,X}\dot{\phi}^{3}\right]\alpha=2M_{\rm P}^{2}\dot{\zeta}-n_{\rm m}\delta l_{\rm m}-n_{\rm c}\delta l_{\rm c}\,.$ (25)
@@ VIII.1 $H(z)$ data
VIII.2 Redshift Space Distortion data
@@
The set of orthogonal functions (grammar) that we have used in our GA analysis are $x$ and $x^{x}$. We have chosen the crossover rate and the mutation rate as $0.75$ and $0.25$ respectively, with $1500$ generations. The algorithm will be terminated once the $\chi^{2}$ does not improve for a few hundred generations. The resulting curve is given in the figure 5. We also plotted the $f\sigma_{8}$ predicted by the model choosing one particular value of $\gamma=0.01$, which is the new parameter in the model. It is clear that the value of $f\sigma_{8}$ in the redshift close to today deviates from the $1\sigma$ predicted by GA, using the redshift space distortion (RSD) data compilation as given in the table 2.
@@
In fact, we have only analysed this model with a particular value of the new parameter. It would be interesting to know if there are other values for the $\gamma$, which can fit with the data. To study this we need to do a complete parameter estimation using Markov chain Monte Carlo (MCMC) methods. This will also allow us to confront the model with other data sets like cosmic microwave background and other large scale structure data, which will be done in a future work.
@@ IX Conclusion
There are several future directions that are interesting to pursue, gaining intuitions from the current work. One direction is to generalize the concrete model to a general tracker solution with $\dot{\phi}H^{p}=\text{constant}$ as explored in [49], and following similar analysis and see if the new parameter $p$ can help in reducing the $f\sigma_{8}$ in to $1\sigma$ region of the genetic algorithm pipeline, using the same data. Another direction is to generalise the interacting model itself as studied in the context of Horndeski theory [79]. Then it would be interesting to see the behavior of the $f\sigma_{8}$. Even though, we did not see any additional propagating modes on the perturbation around the FLRW background, it would be interesting to see if any spurious modes are present in this model. One approach is to do a perturbation analysis on the Bianchi type I spacetime as done in [80, 81, 82]. Another obvious direction is to study the constraints of this model by implementing the Einstein - Boltzmann code and confront with the cosmological data. It is also interesting to study the constraints of this new model with integrated Sachs-Wolfe effect and the large scale structure cross correlation data [83, 84], and study, if the negative cross-correlation is still existing in this model [37, 38]. Another interesting question is the consequence of the model in the non-linear region, where the velocity of the dark matter is significant, and how the screening mechanism is functioning in this model. These studies need an extensive attention, which will be carried out in other works.
@@ Appendix B Behaviour of $G_{cc}$ and $G_{bb}$ along the redshift
In this appendix we show the behaviour of the $G_{cc}/G_{N}$ and $G_{bb}/G_{N}$ along the variable $\mathcal{N}=\ln(a)$, Fig. 6. As we can see that there is enhancement in the ratio between $G_{cc}(G_{bb})$ and $G_{N}$ in the dark energy domination era.